# Observing Double White Dwarfs with the Lunar GW Antenna

### v1

@@ 3.6.1 Error budget on extragalactic population
  • Incompleteness of the HyperLeda catalog: as shown in fig. 4, it is not possible to identify the galaxies that are behind the galactic plane; this will cause a systematic underestimation of the sources that will be visible to LGWA even if obscured by the MW disk for electromagnetic observations. This systematic error does not exceed the $25\%$ of the total population, correspondent to the Zone of Avoidance caused by the obstruction of the MW (Kraan-Korteweg & Lahav, 2000). The bias due to incompleteness would be between $20\%$ and $25\%$ if not properly corrected.
  • Errors in the rate estimation due to uncertainty in the parameters reported in tab. 3: the SNuB coefficients present a moderate errors of approximately $5\%$, the corrective parameter RSSB around $60\%-90\%$. This is due to the limited amount of SN Ia used to calibrate the parameters. The induced statistical error is $\approx 16\%$, obtained by Monte Carlo method in a 30 Mpc radius. We estimate the potential systematic error to be $\approx 5\%$ by comparing the two possible rate calculation procedures (rates obtained preferably from either color or morphological type, see fig. 5)
@@
4 Population analysis
The analysis of the gravitational-wave signal emitted by the population generated in the last section is performed by the codes GWFish (Dupletsa et al., 2023), used as a preferred tool to characterize an accurate response, and LEGWORK (Wagg et al., 2022a, b), used as an alternative to estimate the SNR in the case that the system is not suitable for the GWFish elaboration (spiraling, monochromatic sources).
@@ 4.1 GWFish
This tool enables accurate simulations of the response of a GW detector to a signal characterized by the parameters simulated in sec. 3. A brief explanation of the Fisher-matrix formalism implemented by GWFish is provided in Appendix B. At present, the GWFish code can provide excellent simulations of the detector’s response when the source is relatively close to the merging, but cannot simulate a nearly stationary low-mass source far from the merger. This does not imply that LGWA cannot detect these systems, which can become visible after 10 years of signal integration. However, this is a limit for the analysis of the synthetic population that has been generated in sec. 3, since only a few of the galactic DWDs are suitable for the GWFish analysis.
For this reason, in the following is presented an analysis of the LGWA sensitivity obtained by GWFish referred only to the merging events. Thus, the max_frequency_cutoff parameter is set to be the Roche overflow frequency that characterize the merging (black marks in fig. 3) for a conservative estimation, or the contact frequency for a more realistic but possibly slightly over-optimistic scenario. In addition, only super-Chandrasekhar binaries will be considered. Note that during the observational period the probability of observing a galactic SN Ia is negligible, while the expected SN Ia rate for extragalactic population is such that some merging events are expected within the mission lifetime. However, the super-Chandrasekhar DWDs are only a sub-population of the entire set of (potentially) observable DWDs which can merge during the observational period.
On the contrary, for the extragalactic population we expect to be able to detect only merging events with $\text{min}(m_{1},m_{2})>0.6\,M_{\odot}$ as seen in sec. A. This population requires an accurate model, which can be provided by GWFish. However, also for this population some systems cannot be processed and are thus simply discarded.
@@ 4.2 LEGWORK
The great majority of the spiraling DWDs inside the MW, with a frequency cutoff due to the end of the observation period and not due to the merging of the system, cannot be elaborated by GWFish due to excessively small frequency evolution. We compute an estimation of the SNR for these objects with the code LEGWORK, originally developed for LISA and thus already implementing the SNR calculation for monochromatic sources. The LGWA noise PSD provided in the GWFish resources is used to implement the LGWA response into the code. Note that the SNR calculated in this way is only a rough approximation, since the Doppler shift due to the detector proper motion, the source position in the sky and the inclination are not considered for every specific object. The result is a mean SNR, averaged over this parameter subspace, that is anyway useful to estimate global properties of the population.
4.3 SNR of the MW population
Figure 6: Cumulative distribution of the SNR for the three MW components (colored) and total (black). The plot indicates the cumulative distribution, namely for each SNR threshold $\rho$ (x-axis) the counts (y-axis) provide the number of objects with the SNR larger than the threshold $\rho$. The power-law fits are also reported explicitly. The gray curve indicates the total population as seen by the pathfinder mission Soundcheck.
The SNR distribution resulting from the LEGWORK elaboration is presented in fig. 6. The plot presents the cumulative distribution of the sources: for every SNR $\rho$, the graph gives the number of sources with SNR ¿ $\rho$. In this way it’s easy to quantify the number of observable objects, once a certain SNR threshold is fixed. This distribution follows a power-law $\Gamma=\alpha\cdot\rho^{\beta}$: the exponents $\beta$ for the three populations are not exactly the same due to different characteristics (mass distributions and spatial distributions), but they are sufficiently similar to allow to fit also the total distribution with a power-law. In addition, the majority of the observable sources comes from the thin disk only. For typical SNR thresholds of 8 and 12 are expected respectively 30 and 17 visible sources (with statistical error given by Poissonian statistics). It’s also shown in gray the SNR estimation for a 10 year observational period for the pathfinder mission Soundcheck. Note that, even if this estimation predicts no detections of spiraling DWDs, a detection of a DWD system is not completely ruled out since the given values are to be considered more as an order-of-magnitude estimation and not a precise prediction. Indeed it’s possible the presence of short-period DWD systems not yet identified in the solar neighborhood (Munday et al., 2025).
The galactic inspiraling population emits in the most sensible band of LISA, and thus it will not be a prerogative of LGWA only. In fact, repeating the same estimation using the LISA PSD results is a SNR greater by 3 orders of magnitude on average.

### v2

@@ 3.7 Error budget
  • Initial conditions and parametrization: the population is generated starting from a set of fiducial distributions and parameters. Note however that the total abundance of objects is set independently by the SN Ia rate: for this reason, changes of the initial parameters do not propagate on the global number of super-Chandrasekhar sources, which represent the main target population for LGWA (see Sect. 5). The effect of slightly different parametrizations is only a subdominant rearrangement of the mass distribution. We estimate the error contribution from the initial conditions parametrization to the mass and S/N distributions to be $\lesssim 3\%$ for the super-Chandrasekhar and $\lesssim 7\%$ for the sub-Chandrasekhar subpopulations. Such effects are analyzed and discussed separately in App. C.
  • Incompleteness of the HyperLeda catalog: as shown in fig. 3, it is not possible to identify the galaxies that are behind the Galactic plane; this will cause a systematic underestimation of the sources that will be visible to LGWA even if obscured by the MW disk for electromagnetic observations. This systematic error does not exceed the $25\%$ of the total population, correspondent to the Zone of Avoidance caused by the obstruction of the MW (Kraan-Korteweg & Lahav, 2000). The bias due to incompleteness would be between $20\%$ and $25\%$ if not properly corrected. Although we always apply a correction, a potential systematic error of $\approx 5\%$ is also taken into account.
  • Errors in the rate estimation due to uncertainty in the parameters reported in tab. 3: the SNuB coefficients present a moderate errors of approximately $5\%$, the corrective parameter RSSB around $60\%-90\%$. This is due to the limited amount of SN Ia used to calibrate the parameters. The induced statistical error is $\approx 16\%$, obtained by Monte Carlo method in a 30 Mpc radius. We estimate the potential systematic error to be $\approx 5\%$ by comparing the two possible rate calculation procedures (rates obtained preferably from either color or morphological type, see fig. 11). Correspondingly, the Galactic population inherits from the adopted SN Ia rate an uncertainty of $\approx 20\%$.
@@
[…]
4 Gravitational-wave analysis
The analysis of the gravitational-wave signal emitted by the populations generated in the last section is performed by the codes GWFish (Dupletsa et al., 2023), used as a preferred tool to characterize an accurate response, and LEGWORK (Wagg et al., 2022a, b), used as an alternative to estimate the S/N in the case that the system is not suitable for the GWFish elaboration. We distinguish between “inspiraling” sources, for which the frequency cutoff is due to the end of the observational period, and “merging” sources, for which the frequency cutoff is due to the merging of the binary. In particular we provide an analysis of the entire MW population, which consists of inspiraling sources, in Sect. 5.1; the MW merging population is treated in Sect. 5.2. Regarding the extragalactic population, in Sect. 5.3 is presented the analysis of the only super-Chandrasekhar merging population, since as demonstrated in Appendix A these are the only observable systems outside the MW. In the following we present an introduction to the used codes.
@@ 4.1 GWFish
This tool enables accurate simulations of the response of a GW detector to a signal characterized by the parameters simulated in Sect. 3, providing an estimation of the errors on the inferred values of the parameters using the Fisher-matrix formalism (see App. D). At present, the GWFish code can provide excellent simulations of the detector’s response when the source is relatively close to the merging, but cannot simulate a nearly stationary low-mass source far from the merger. This does not imply that LGWA cannot detect these systems, which can become visible after 10 years of signal integration. For this reason, in the following is presented an analysis of the LGWA sensitivity obtained by GWFish referred only to the merging events. Thus, the max_frequency_cutoff parameter is set to be the Roche overflow frequency that characterize the merging (black marks in fig. 10) for a conservative estimation, or the contact frequency for a more realistic but possibly slightly over-optimistic scenario. Note that during the observational period the probability of observing a Galactic SN Ia is negligible, while the expected SN Ia rate for extragalactic population is such that some merging events are expected within the mission lifetime. GWFish is thus used to precisely characterize merging extragalactic systems and merging super-Chandrasekhar systems.
@@ 4.2 LEGWORK
We compute an estimation of the S/N for monochromatic MW binaries with the code LEGWORK, originally developed for LISA and thus already implementing the S/N calculation for monochromatic sources. The LGWA noise PSD provided in the GWFish resources is used to implement the LGWA response into the code. Note that the S/N calculated in this way is only a rough approximation, since the Doppler shift due to the detector proper motion, the source position in the sky and the inclination are not considered for every specific object. The result is a mean S/N, averaged over this parameter subspace, that is anyway useful to estimate global properties of the population.
5 Results
5.1 S/N of the MW inspiraling population
Figure 4: Cumulative distribution of the S/N for the three MW components (colored) and total (black) over ten years of LGWA observation. The plot indicates the cumulative distribution, namely for each S/N threshold $\rho$ (x-axis) the counts (y-axis) provide the number of objects with the S/N larger than the threshold $\rho$. The power-law fits are also reported explicitly.
The S/N distribution resulting from the LEGWORK elaboration of the entire (super- and sub- Chandrasekhar) MW inspiraling population is presented in fig. 4. The plot presents the cumulative distribution of the sources: for every $\rho$, the graph gives the number of sources with S/N ¿ $\rho$. This directly quantifies the number of observable objects, once a certain S/N threshold is fixed. This distribution follows a power-law $\Gamma=\alpha\cdot\rho^{\beta}$: the exponents $\beta$ for the three populations are not exactly the same due to different characteristics (mass and spatial distributions), but they are sufficiently similar to allow to fit also the total distribution with a power-law. In addition, the majority of the observable sources comes from the thin disk only. The $\alpha$ coefficients correspond to the number of sources with S/N ¿ 1. For typical S/N thresholds of 8 and 12, over an observing period of ten years $30\pm 5\text{(stat)}\pm 6\text{(sys)}$ and $17\pm 4\text{(stat)}\pm 4\text{(sys)}$ visible sources are expected, respectively.
The Galactic inspiraling population emits in the most sensitive band of LISA, and thus it will not be a prerogative of LGWA only. Repeating the same estimation using the LISA PSD results in a systematically larger S/N. The simultaneous detection of common sources can however be relevant both for scientific and instrumental purposes, such as detector calibration.